GlueX Start Counter Meeting, June 25, 2020
- Review minutes from the last meeting
- Calibration and Efficiency
- Dark Rate Studies
- Action Item Review
Present: Eugene Chudakov, Rupesh Dotel, Tolga Erbora, Mark Ito (chair), Simon Taylor, Beni Zihlmann
There is a recording of this meeting on the BlueJeans site. Use your JLab credentials to gain access.
Review of minutes from the last meeting
We went over the minutes from May 28.
Time Resolution Studies
Rupesh showed plots of time resolution for all thirty counters for a run late in the last run period, run 71840. Find his plots here (starting with slide 3). Recall that last time he showed evidence for worsening of the resolution as a function of time as the run progressed when using calibration constants obtained early in the run. This time constants were obtained for run 71840 and applied to the same run. For this run, results were not much different than those with the earlier-in-the-run constants. He also tried a run even later in the run, 72246, which showed slightly better resolution. To see if there was anything special about run 71840, he looked other runs in the same run-number neighborhood. Those were all roughly the same as for 71840.
Eugene asked about the source of the counter-to-counter variation in resolution, noting that the pattern repeats for all runs shown. Beni and Simon attribute that to the rate variation from counter-to-counter due to a slight tilt of the assembly and the concentration of the rate at the nose of the counters where small variations in distance from the beam make a big difference in rate.
After the last meeting Sean Dobbs reported via email that an initial SC gain calibration was done by Simon back in 2015, and that has been what we are using so far. Beni also pointed out that the Start Counter NIM paper had dE/dx plots both before and after "corrections." Rupesh and Beni will look at some recent monitoring plots to see if they give an indication about the current adequacy of the calibration.
Improving Speed of Performing Calibrations
Rupesh consulted Beni after the last meeting and found that he could reduce the number of files analyzed for a calibration pass from the full run (over 200 raw data files) to 25 and still get reliable results. That turns a month into a few days.
Beni raised the issue of how we might approach replacement of the SiPMs if we conclude, either now or in the future, that radiation damage is too severe to continue to use the current units. We need to talk to Fernando about whether it would be easier to replace the SiPM on the current boards or just make new boards with new SiPMs. We would have to research the availability of the current units. Also, we would need a plan to do the replacement mechanically including locating any necessary support jigs or structures.
Preparations for the Upcoming Run
Beni proposes that we leave the start counter alone until after the Fourth of July. After turn-on, he wants to take dark-current measurements (non-sparsified raw mode data) before beam arrives in the Hall. He recalled for us that that attempt before the previous run was mis-configured, DAQ-wise, and the resulting data unusable.
Action Item Review
- Document calibration procedure (Rupesh)
- Reproduce Beni's dark current results. (Tolga)
- Check status of the gain calibrations using monitoring histograms. (Rupesh, Beni)
- Revisit Yi's study of SiPM annealing. (Mark, Beni)
- Ask about the current recommendation from SciComp on copying files to the local disk from cache before analyzing them. (Mark)
- Ask Fernando about strategy for SiPM replacement (Mark, Beni)